OpenStudy (darkbluechocobo):

Apply the mathematical relationship between energy and amplitude to describe the effect on the energy of a wave when the amplitude is tripled.

OpenStudy (darkbluechocobo):

@Data_LG2

OpenStudy (anonymous):

E=kA^2 am i right? so it means...\[\sf E\ \alpha\ A^2\] if we triple the amplitude.. E'= (3A)^2 E'= 9A^2 E' = 9E it means the energy will be 9 times the initial energy

OpenStudy (darkbluechocobo):

Sorry Im back

OpenStudy (anonymous):

no worries (: i have to go though... if you have more questions just tag wio-good at physics :D see ya!

OpenStudy (darkbluechocobo):

Alright will do :p